OpenStudy (anonymous):

Which of the following properly shows the properties of negative numbers? -4 • 9 = 36 3 • (-5) = 15 -4 • (-9) = -36 4 • (-9) = -36

OpenStudy (anonymous):

A good way to look at it is this way. Whenever you multiply/divide a positive number with a negative number you will receive a negative number. When you multiply/divide a positive and positive number you get a positive. When you multiply/divide a negative by a negative you get a positive number.

OpenStudy (anonymous):

what?! -_^

OpenStudy (anonymous):

+ x + = + - x - = + + x - = -

OpenStudy (anonymous):

ooooo ok i got that! what does that have to do with the problem though

OpenStudy (anonymous):

Then look at your option of answers. Which is correct?

OpenStudy (anonymous):

4 * (-9) = -36

OpenStudy (anonymous):

The first is wrong because + x - = - the second one which is -4 x -9 = -36 is wrong because - x - = +

OpenStudy (anonymous):

The third one 3 x -5 = 15 is wrong because + x - = -

OpenStudy (anonymous):

so i was rite?!?!?

OpenStudy (anonymous):

The answer has to be the 4th one because + x - = - ; 4x-9= -36

OpenStudy (anonymous):

I don't know which answer you picked, so I guess.....?

OpenStudy (anonymous):

Yes sorry didn't see your answer srry x_x

OpenStudy (anonymous):

yeah!!!!!!! :) :) :) :) :) 02~01~2013 0 4 * (-9) = -36

OpenStudy (anonymous):

lol its fine... thanks for all the help!

OpenStudy (anonymous):

You're welcome. (: